# Nonlinear Permuted Granger Causality

### v1

@@ 3.1 Asymptotic Properties
We examine the asymptotic behavior of the estimates and established testing framework under the listed conditions in this section and the supplement. Define the underlying variation parameter for permutation $m$ as $\vartheta_{m}$, and the estimate $\hat{\vartheta}_{m}$ as in Equation 6. We assume a fixed test set size $T_{k}$ and allow the training set $T_{-k}=T-T_{k}$ and number of folds $K$ to grow as $T\rightarrow\infty$.
Theorem 1.
Under the conditions listed in Section 3 and the supplement, with Condition 2 modified such that $\tau_{\omega}^{2}=0$ (i.e., we achieve error free observation of the data), for all $\varepsilon>0$,
@@
Alternatively, under the conditions listed in Section 3 and the supplement, for all $\varepsilon>0$,
@@
Theorem 2.
Under the conditions listed in Section 3 and the supplement, the estimate for the variation parameter $\hat{\vartheta}_{m}$ admits a Central Limit Theorem with respect to the number of observations $\varphi$ in $\Omega_{\text{obs}}$.
@@
As a direct result of Theorem 2, $\hat{\vartheta}_{m}$ is a consistent estimate for $\vartheta_{m}$ when $\tau_{\omega}^{2}>0$ as $\varphi$ tends to infinity along with the size of the training and test sets. Next, we examine the test statistic $\hat{Q}_{M}$ under the null and alternative hypotheses.
@@ 3.1.1 Under the Null Hypothesis
Under the null hypothesis, the underlying variation parameter is constant for every possible permutation, $\vartheta_{1}=\cdots=\vartheta_{T}!$. We define the quantile estimate $\hat{Q}_{M}$ as in Equation 9, and can establish the limiting uniform distribution from the result of Theorem 1.
Theorem 3.
Under the null hypothesis and the conditions listed in Section 3 and the supplement,
@@ 3.1.2 Under the Alternative Hypothesis
We define $Q$ as the true quantile for parameter $\vartheta_{1}$ over all $T!$ possible permutations. In the limit, the quantile estimate defined in Equation 9 converges in probability to $Q$.
Theorem 4.
Under the alternative hypothesis and the conditions listed in Section 3 and the supplement, with Condition 2 modified such that $\tau_{\omega}^{2}=0$ (i.e., we achieve error free observation of the data), for all $\varepsilon>0$,
@@
Similarly, under the alternative hypothesis and the conditions listed in Section 3 and the supplement, for all $\varepsilon>0$,
@@ 4 Simulation Study
We evaluate the ability of the permutation-based methodology of NPGC (similar to the decision rule formulation of TCDF) to detect the presence of functional connectivity and control for false positive results with a consistent single-layer cMLP nonlinear transformation framework. All methods use featurization dimension $N=100$. Three of the five comparison methods use in-sample methodology and variants of the lasso penalty within cMLP models (Tank et al., 2022; Tibshirani, 1996). The other two methods employ random initialization like NPGC, and use in-sample testing and decision rules from comparison of restricted and unrestricted models or the substitution of random Gaussian noise.
Multiple testing on bivariate pairs is not examined; we consider a group decision process where collectively the covariate set $\mathbf{X}$ is (or is not) Granger causal for the response $\mathbf{Y}$. Ability to detect a Granger causal result will almost certainly improve with a more sophisticated nonlinear structure; we demonstrate the relative differences between methods at a baseline level. A full list of included methods is given in Table 1.
(i) Nonlinear Permuted Granger Causaltiy (NPGC)
@@
The lasso type objectives of cMLP methods (ii) - (iv) penalize nonzero rows of the matrix $\mathbf{W}^{1}$ in the formulation of Equation 1. The corresponding null and alternative hypotheses are adjusted to those shown in Equation 18, and a Granger causal connection is present for a covariate set if any corresponding rows of $\mathbf{W}^{1}$ contain nonzero components (Tank et al., 2022).
@@
A group lasso penalty is applied to each variable $i$ in the model matrix $\mathbf{W}^{1}$ corresponding to the covariate set $\mathbf{X}$ over all time lags $t,\ldots,t-\gamma$ (Yuan and Lin, 2006). As in Tank et al. (2022), we define $\mathbf{W}^{1}_{i}=\left[\mathbf{W}^{1}_{i,t}\;\cdots\;\mathbf{W}^{1}_{i,t-\gamma}\right]$, with each entry as the weights of the model matrix row for variable $i$ and lagged data up to $t-\gamma$. Penalties for the cMLP methods take the general form $\lambda\sum_{i=1}^{p}\beta\left(\mathbf{W}^{1}_{i}\right)$, with specific $\beta$ for each defined below.
$\beta_{GL}\left(\mathbf{W}^{1}_{i}\right)$ $=\left\|\mathbf{W}^{1}_{i}\right\|_{F}$
$\beta_{GSGL}\left(\mathbf{W}^{1}_{i}\right)$ $=\left\|\mathbf{W}^{1}_{i}\right\|_{F}+\sum_{k=0}^{\gamma}\left\|\mathbf{W}^{1}_{i,t-k}\right\|_{2}$
$\beta_{H}\left(\mathbf{W}^{1}_{i}\right)$ $=\sum_{k=0}^{\gamma}\left\|\left[\mathbf{W}^{1}_{i,t-k}\;\cdots\;\mathbf{W}^{1}_{i,t-\gamma}\right]\right\|_{F}$ (19)
The group sparse group lasso penalty combines sparsity of included variables and their lagged values like in Simon et al. (2013), and the novel hierarchical penalty of Tank et al. (2022) retains information about the natural ordering of the variables, encouraging solutions where for some lag $^{*}, k>k$$k^{*}$ implies $\mathbf{W}^{1}_{i,t-k}=\mathbf{0}$. We select a regularization parameter of $\lambda=0.5$ based on a cross-validation like trial and error approach. Conclusions drawn from these methods can vary greatly depending on the chosen $\lambda$; smaller values retain all matrix entries and larger values penalize the matrix to zero.
We formulate two naive methods for comparison to NPGC. The in-sample restricted and unrestricted comparison examines the ratio of model residuals $\hat{\sigma}^{2}_{\text{Res}}/\hat{\sigma}^{2}_{\text{Unres}}$ in a randomly initialized cMLP model. The in-sample Gaussian noise substitution is methodologically similar, but instead substitutes Gaussian noise in place of the covariate set $\mathbf{X}$ to examine $\hat{\sigma}^{2}_{\text{Noise}}/\hat{\sigma}^{2}_{\text{Unres}}$. The three model errors are shown in Equation 20, where $\mathbf{E}\sim\mathcal{MN}_{T\times p}\left(\mathbf{0},\mathbf{I},\mathbf{I}\right)$, and the corresponding $\hat{\sigma}^{2}$ shown in Equation 21.
$\mathbf{U}_{\text{Unres},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{X}\right]\mathbf{W}_{r}\right)\mathbf{B}_{\text{Res},r}$
$\mathbf{U}_{\text{Res},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{0}\right]\mathbf{W}_{r}\right)\mathbf{B}_{\text{Res},r}$
$\mathbf{U}_{\text{Noise},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{E}\right]\mathbf{W}_{r}\right)\mathbf{B}_{\text{Noise},r}$ (20)
@@
The individual terms in Equation 21 follow chi-square distributions with degrees of freedom TNT-N, and each ratio for an individual initialization is distributed $\mathcal{F}_{T-N,T-N}$. Over all initializations $r=1,\ldots,\mathscr{R}$ the sum of a large number of these independent (via Condition 3) statistics is approximately normal. As $\mathscr{R}\rightarrow\infty$ under the null hypothesis,
@@
We formulate a naive decision rule with this limiting distribution and a large number ($\mathscr{R}=1000$) of random initializations.
The NPGC methodology is performed with $M=400$ permutations, $K=5$ cross validation folds, and $\mathscr{R}=50$ random initializations. We evaluate the methods based on correctly flagging a Granger causal result when direct functional dependence is present, or correctly labelling a non-causal result when it is absent, leading to the potential outcomes in Table 2. For the NPGC and naive ratio methods, we flag a result if the quantile estimate of the associated statistic under the null hypotheses is under a specified level $\alpha$. The lasso-type methods lack the direct translation to a traditional hypothesis testing framework.
Result $\backslash$ Truth Granger causal ($\text{GC}=1$) Not causal ($\text{GC}=0$)
@@
Table 2: Proportion of potential outcomes for the decision $\mathcal{D}$ of each of the methods shown in Table 1. The set $\mathcal{S}$ represents the space of all simulations (both $\text{GC}=1$ and $\text{GC}=0$) within a given setting, $N_{1}=\sum_{\mathcal{S}}\mathbf{1}\left\{\text{GC}=1\right\}$, and $N_{0}=\sum_{\mathcal{S}}\mathbf{1}\left\{\text{GC}=0\right\}$.
@@ 4.1 Simulation Settings
We test NPGC and the comparator methods in Table 1 on two nonlinear processes: Lorenz-96 models of Karimi and Paul (2010), and threshold autoregressive models (TAR) introduced by Tong (1978). The $p$-dimensional Lorenz-96 model ($p\geq 4$) is governed by the continuous differential equation
@@
with $i=1,\ldots,p$ and boundary series $x_{-1,t}=x_{p-1,t}$, $x_{0,t}=x_{p,t}$, and $x_{p+1,t}=x_{1,t}$. $F$ is a forcing constant we generate as $F\sim\mathcal{U}(5,20)$ with higher values introducing a larger degree of nonlinear, chaotic behavior. For data generation, we use a sampling rate of $\Delta t=0.05$ and allow a burn-in period of 500 time steps.
The TAR(2) model of Tong (1978) is governed by a similar skeleton to a vector autoregressive model, but allows for changes in parameters based on the value of a threshold variable (Tong and Lim, 1980).
@@
We define the threshold on the values in $\mathbf{x}_{t-2}$ with regime $k=1$ corresponding to the case when $\sum_{i=1}^{p}x_{i,t-2}\leq 0$ and $k=2$ to $\sum_{i=1}^{p}x_{i,t-2}>0$. For the TAR(2) process, $\bm{\varepsilon}^{(k)}_{t}\sim\mathcal{N}_{p}\left(\mathbf{0},\sigma_{k}^{2}\mathbf{I}\right)$ where $\sigma_{1}=0.5$ and $\sigma_{2}=0.2$. Each $\mathbf{A}^{(k)}_{1},\mathbf{A}^{(k)}_{2}\in\mathbb{R}^{p\times p}$ has elements $\mathbf{A}^{(k)}=(a^{(k)}_{ij})\sim\mathcal{U}(-0.5,0.5)$ that are thresholded to zero if $|a^{(k)}_{ij}|<0.1$, and their spectral radii are at most 0.8 to ensure stationary data generation. From the initial zero state, we allow for a burn-in period of 500 time steps as in the Lorenz-96 data generation procedure.
We generate two groups of dependent data observations with $p=6$ (one from each process), containing $\mathbf{x}_{i}$ for $i=1,\ldots,12$. We label $i=1,\ldots,6$ as the Lorenz-96 process and $i=7,\ldots,12$ as the TAR(2) process. The samples are of length $T=250$, 500500, and $1000$ after burn-in and truncation for lag selection to include in the model, $\gamma=3$. We select one index to serve as the response, three to serve as the set of additional variables $\mathbf{Z}$ and three for the covariate set $\mathbf{X}$ depending on a Granger causal designation. A dataset is designated “Granger causal” if at least one series in $\mathbf{X}$ has a direct influence on the selected response (i.e., one series in $\mathbf{X}$ is chosen from the same generating process). We simulate 200 datasets for each of the settings shown in Table 3 for a total of 4800 trials.
Response Variable $T$ # Causal in $\mathbf{X}$ (GC) # Causal in $\mathbf{Z}$
@@
Table 3: Settings for simulation and selection of variables. A variable is labelled causal if from the same generating process as the response $\mathbf{Y}$. Variables are randomly selected from within their causal or non-causal groups.
@@ 4.2 Simulation Results
We present simulation results split by Granger causal designation and process of the chosen response variable. For each, we examine the control for false positive results and the ability to detect a Granger causal connection. We do not examine AUROC as it does not implicate a decision rule a priori. Tables 4(a) and 4(b) list the proportion of correctly identified causal relationships $\rho_{1}$ and Figures 1 and 2 the observed Type 1 Error by specified level of test. The results for the finer grid of designations from Table 3 are given in the supplement.
@@
(a) TAR(2) response
@@
(b) Lorenz-96 response
Table 4: Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2). Results are presented by the fine grid of designations in Table 3 in the supplementary material.
The NPGC method succeeds in identifying many cases of Granger causal influence in both processes, and as expected, this ability approaches the upper limit of one as the number of time points increases. The cMLP methods provide spotty results, and these are greatly influenced by the selected parameter $\lambda$. Without guidance for a specific penalty selection, the global decision of “Granger causal” or “non-causal” is uncertain. The naive restricted and unrestricted model method fails to identify many Granger causal pairings in the Lorenz-96 trials. The Gaussian noise substitution method correctly identifies nearly all Granger causal pairs, but this comes at the cost of uncontrolled Type 1 error (see Figures 1 and 2).
Figure 1: Type 1 Error control by reference quantile $\alpha$ for TAR(2) simulations. Methods that adhere to a chosen level of test will lie closer to the uniform cdf (gray) included for reference.
Figure 2: Type 1 Error control by reference quantile $\alpha$ for Lorenz-96 simulations. Methods that adhere to a chosen level of test will lie closer to the uniform cdf (gray) included for reference.
Under the null hypothesis, we demonstrate moderate adherence to the asymptotic properties shown in Theorem 3 by presenting the observed Type 1 Error control by generic reference quantile $\alpha$. The NPGC permutation methodology tracks relatively close to the included uniform cdfs in Figure 1 for TAR(2) simulation settings, and exhibits mild to moderate deviations for Lorenz-96 settings in Figure 2. We can likely attribute these deviations to the complicated nature of the chaotic Lorenz system; we do not see this same pattern in other simulations when both groups are generated by a TAR(2) process (see Figure 7 in the supplement).
The cMLP lasso methods are included as horizontal lines in these plots; they do not have a straightforward extension for a specific level of test, other than computationally expensive iteration over several penalties $\lambda$. The naive methods do not adhere to their theoretical Type 1 error control, but approaches in the noise substitution realm appear promising if this issue can be resolved.

### v2

@@ 3.1 Asymptotic Properties
The asymptotic behavior of the estimates and established testing framework is examined under the listed conditions above and in the supplement. Define the underlying variation parameter for permutation $m$ as $\vartheta_{m}$, and the estimate $\hat{\vartheta}_{m}$ as in Equation 13. Assume a fixed test set size $T_{k}$ and allow the training set $T_{-k}=T-T_{k}$ and number of folds $K$ to grow as $T\rightarrow\infty$.
Theorem 3.1.
Under the listed conditions, with Condition 3 modified such that $\tau_{\omega}^{2}=0$ (i.e., error free observation of the data), for all $\varepsilon>0$,
@@
Alternatively, under the listed conditions, for all $\varepsilon>0$,
@@
Theorem 3.2.
Under the listed conditions, the estimate for the variation parameter $\hat{\vartheta}_{m}$ admits a Central Limit Theorem with respect to the number of observations $\varphi$ in $\Omega_{\text{obs}}$.
@@
As a direct result of Theorem 3.2, $\hat{\vartheta}_{m}$ is a consistent estimate for $\vartheta_{m}$ when $\tau_{\omega}^{2}>0$ as $\varphi$ tends to infinity along with the size of the training and test sets.
@@ 3.1.1 Under the Null Hypothesis
Under the null hypothesis, the underlying variation parameter is constant for every possible permutation, $\vartheta_{1}=\cdots=\vartheta_{T}!$. Define the quantile estimate $\hat{Q}_{M}$ as in Equation 16, and establish the limiting uniform distribution from the result of Theorem 3.1.
Theorem 3.3.
Under the null hypothesis and the listed conditions,
@@ 3.1.2 Under the Alternative Hypothesis
Define $Q$ as the true quantile for parameter $\vartheta_{1}$ over all $T!$ possible permutations. In the limit, the quantile estimate defined in Equation 16 converges in probability to $Q$.
Theorem 3.4.
Under the alternative hypothesis and the listed conditions, with Condition 3 modified such that $\tau_{\omega}^{2}=0$ (i.e., error free observation of the data), for all $\varepsilon>0$,
@@
Similarly, under the alternative hypothesis and the listed conditions, for all $\varepsilon>0$,
@@ 4 Simulation Study
The ability of the permutation-based methodology of NPGC (similar to the decision rule formulation of TCDF) to detect the presence of functional connectivity and control for false positive results is evaluated with a consistent FNN framework (like the cMLP nonlinear transformation). All methods featurize the data to dimension $N=100$. Three of the five comparison methods use in-sample methodology and variants of the lasso penalty within cMLP models (Tank et al., 2022; Tibshirani, 1996). The other two employ random FNN generation like NPGC, and in-sample testing from comparison of restricted and unrestricted models or the substitution of random Gaussian noise.
A group decision process is considered where the covariate set $\mathbf{X}$ is (or is not) collectively Granger causal for the response $\mathbf{Y}$. Ability to detect a Granger causal result will almost certainly improve with a more sophisticated nonlinear structure; relative differences between methods are demonstrated at a baseline level. A full list of included methods is given in Table 1.
(i) Nonlinear Permuted Granger Causality (NPGC)
@@
The lasso type objectives of cMLP methods (ii) - (iv) penalize nonzero rows of the matrix $\mathbf{W}^{0}$ in the formulation of Equation 5 (or $\mathbf{W}$ in Equation 10). The corresponding null and alternative hypotheses are adjusted to those shown in Equation 29, and a Granger causal connection is present for a covariate set if any corresponding rows of $\mathbf{W}^{0}$ contain nonzero components (Tank et al., 2022).
@@
A group lasso penalty is applied to each variable $j$ in the model matrix $\mathbf{W}^{0}$ corresponding to the covariate set $\mathbf{X}$ over all time lags $t,\ldots,t-\gamma$ (Yuan and Lin, 2006). As in Tank et al. (2022), define $\mathbf{W}^{0}_{j}=\left[\mathbf{W}^{0}_{j,t}\;\cdots\;\mathbf{W}^{0}_{j,t-\gamma}\right]$, with each entry as the weights of the model matrix row for variable $j$ and lagged data up to $t-\gamma$. Penalties for the cMLP methods take the general form $\lambda\sum_{j=1}^{p}\beta\left(\mathbf{W}^{0}_{j}\right)$, with specific $\beta$ for each defined below.
$\beta_{GL}\left(\mathbf{W}^{0}_{j}\right)$ $=\left\|\mathbf{W}^{0}_{j}\right\|_{F}$ (30)
$\beta_{GSGL}\left(\mathbf{W}^{0}_{j}\right)$ $=\left\|\mathbf{W}^{0}_{j}\right\|_{F}+\sum_{k=0}^{\gamma}\left\|\mathbf{W}^{0}_{j,t-k}\right\|_{2}$ (31)
$\beta_{H}\left(\mathbf{W}^{0}_{j}\right)$ $=\sum_{k=0}^{\gamma}\left\|\left[\mathbf{W}^{0}_{j,t-k}\;\cdots\;\mathbf{W}^{0}_{j,t-\gamma}\right]\right\|_{F}$ (32)
The group sparse group lasso penalty combines sparsity of included variables and their lagged values like in Simon et al. (2013), and the novel hierarchical penalty of Tank et al. (2022) retains information about the natural ordering of the variables, encouraging solutions where for some lag $^{*}, k>k$$k^{*}$ implies $\mathbf{W}^{0}_{j,t-k}=\mathbf{0}$. A regularization parameter of $\lambda=0.5$ is chosen based on a cross-validation like trial and error approach. Conclusions drawn from these methods can vary greatly depending on the chosen $\lambda$; smaller values retain all matrix entries and larger values penalize the matrix to zero. Effects of varying $\lambda$ are shown for the application in Section 5.
Two additional naive methods are included for comparison to NPGC. The in-sample restricted and unrestricted method examines the ratio of model residuals $\hat{\sigma}^{2}_{\text{Res}}/\hat{\sigma}^{2}_{\text{Unres}}$ in a randomly generated FNN. The in-sample Gaussian noise substitution is methodologically similar, but instead substitutes Gaussian white noise in place of the covariate set $\mathbf{X}$ to examine $\hat{\sigma}^{2}_{\text{Noise}}/\hat{\sigma}^{2}_{\text{Unres}}$. The three model errors are shown in Equations 33 to 35, where $\mathbf{E}\sim\mathcal{MN}_{T\times p}\left(\mathbf{0},\mathbf{I},\mathbf{I}\right)$, and the corresponding $\hat{\sigma}^{2}$ shown in Equation 36.
$\mathbf{U}_{\text{Unres},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{X}\right]\mathbf{W}_{r}\right)\mathbf{W}^{L}_{\text{Unres},r}$ (33)
$\mathbf{U}_{\text{Res},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{0}\right]\mathbf{W}_{r}\right)\mathbf{W}^{L}_{\text{Res},r}$ (34)
$\mathbf{U}_{\text{Noise},r}$ $=\mathbf{Y}-\tanh\left(\left[\mathbf{1}\;\mathbf{Y}_{\text{lag}}\;\mathbf{Z}\;\mathbf{E}\right]\mathbf{W}_{r}\right)\mathbf{W}^{L}_{\text{Noise},r}$ (35)
@@
The individual terms in Equation 36 follow chi-square distributions with degrees of freedom TNT-N, and each ratio for an individual generated FNN is distributed $\mathcal{F}_{T-N,T-N}$. The sum of a large number of these independent (via Condition 3) statistics, all random FNNs $r=1,\ldots,\mathscr{R}$, is approximately normal. As $\mathscr{R}\rightarrow\infty$ under the null hypothesis,
@@
This is used to formulate a naive decision rule with a large number ($\mathscr{R}=1000$) of randomly generated FNNs.
The NPGC methodology is performed with $M=400$ permutations, $K=5$ cross validation folds, and $\mathscr{R}=50$ randomly generated FNNs. The methods are successful when correctly flagging a Granger causal result when direct functional dependence is present, or correctly labelling a non-causal result when it is absent, leading to the potential outcomes in Table 2. For the NPGC and naive ratio methods, a result is flagged if the quantile estimate of the associated statistic under the null hypotheses is under a specified level $\alpha$. The lasso-type methods lack the direct translation to a traditional hypothesis testing framework.
Truth
Result Granger causal ($\text{GC}=1$) Not causal ($\text{GC}=0$)
@@
Table 2: Proportion of potential outcomes for the decision $\mathcal{D}$ of each of the methods shown in Table 1. The set $\mathscr{S}$ represents the space of all simulations (both $\text{GC}=1$ and $\text{GC}=0$) within a given setting, $N_{1}=\sum_{\mathscr{S}}\mathbf{1}\left\{\text{GC}=1\right\}$, and $N_{0}=\sum_{\mathscr{S}}\mathbf{1}\left\{\text{GC}=0\right\}$.
@@ 4.1 Simulation Settings
NPGC and the comparator methods in Table 1 are tested on two nonlinear processes: Lorenz-96 models of Karimi and Paul (2010), and TAR models introduced by Tong and Lim (1980). The $p$-dimensional Lorenz-96 model ($p\geq 4$) is governed by the continuous differential equation
@@
with $i=1,\ldots,p$ and boundary series $x_{-1,t}=x_{p-1,t}$, $x_{0,t}=x_{p,t}$, and $x_{p+1,t}=x_{1,t}$. $F$ is a forcing constant generated as $F\sim\text{Uniform}(5,20)$ with higher values introducing a larger degree of nonlinear, chaotic behavior. For data generation, a sampling rate of $\Delta t=0.05$ and a burn-in period of 500 time steps are used.
The TAR(2) model is governed by a similar skeleton to a VAR, but allows for changes in parameters based on the value of a threshold variable (Tong and Lim, 1980).
@@
The threshold is defined on the values in $\mathbf{x}_{t-2}$ with regime $k=1$ corresponding to the case when $\sum_{i=1}^{p}x_{i,t-2}\leq 0$ and $k=2$ to $\sum_{i=1}^{p}x_{i,t-2}>0$. For the TAR(2) process, $\bm{\varepsilon}^{(k)}_{t}\sim\mathcal{N}_{p}\left(\mathbf{0},\sigma_{k}^{2}\mathbf{I}\right)$ where $\sigma_{1}=0.5$ and $\sigma_{2}=0.2$. Each $\mathbf{A}^{(k)}_{1},\mathbf{A}^{(k)}_{2}\in\mathbb{R}^{p\times p}$ has elements $\mathbf{A}^{(k)}=(a^{(k)}_{ij})\sim\text{Uniform}(-0.5,0.5)$ that are thresholded to zero if $|a^{(k)}_{ij}|<0.1$, and their spectral radii are at most 0.8 to ensure stationary data generation. Like the Lorenz-96 data generation procedure, there is a burn-in period of 500 time steps.
Two groups of dependent data observations are generated with $p=6$ (one from each process), containing $\mathbf{x}_{i}$ for $i=1,\ldots,12$. The series $i=1,\ldots,6$ are labelled as the Lorenz-96 process and $i=7,\ldots,12$ as the TAR(2) process. The samples are of length $T=250$, $500$, and $1000$ after burn-in and truncation for lag selection to include in the model, $\gamma=3$. One index is selected to serve as the response, three to serve as the set of additional variables $\mathbf{Z}$ and three for the covariate set $\mathbf{X}$ depending on a Granger causal designation. A dataset is designated “Granger causal” if at least one series in $\mathbf{X}$ has a direct influence on the selected response (i.e., one series in $\mathbf{X}$ is chosen from the same generating process). For each of the settings in Table 3, 200 datasets are generated for a total of 4800 trials.
Response Variable $T$ # Causal in $\mathbf{X}$ # Causal in $\mathbf{Z}$
@@
Table 3: NPGC simulation settings. A variable is labelled causal if from the same generating process as the response $\mathbf{Y}$. Variables are randomly selected from within their causal or non-causal groups.
@@ 4.2 Simulation Results
Simulation results are split by Granger causal designation and process of the chosen response variable. The control for false positive results and the ability to detect a Granger causal connection is examined in each setting. AUROC is not considered as it does not implicate a decision rule a priori. Tables 4 and 5 list the proportion of correctly identified causal relationships $\rho_{1}$ for TAR(2) and Lorenz-96 response variables, respectively. Results for the finer grid of designations from Table 3 are given in the supplement.
@@
Table 4: TAR(2) response compiled simulation results. Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2).
@@
Table 5: Lorenz-96 response compiled simulation results. Proportion of correctly labelled Granger causal outcomes ($\text{GC}=1$, $\rho_{1}$ in Table 2).
Figures 1 and 2 plot the observed Type 1 error by specified level of test, with a target 45 degree line included for reference.
Figure 1: Type 1 error control for TAR(2) simulations. Methods that adhere to a chosen level of test will lie closer to the uniform CDF (gray) included for clarity that indicates the reference quantile $\alpha$.
Figure 2: Type 1 error control for Lorenz-96 simulations. Methods that adhere to a chosen level of test will lie closer to the uniform CDF (gray) included for clarity that indicates the reference quantile $\alpha$.
The NPGC method succeeds in identifying many cases of Granger causal influence in both processes, and as expected, this ability approaches the upper limit as the number of time points increases. The cMLP methods provide spotty results, and these are greatly influenced by the selected parameter $\lambda$. Without guidance for a specific penalty selection, the global decision of “Granger causal” or “non-causal” is uncertain. The naive restricted and unrestricted model method fails to identify many Granger causal pairings in the Lorenz-96 trials. The Gaussian noise substitution method correctly identifies nearly all Granger causal pairs, but this comes at the cost of uncontrolled Type 1 error (see Figures 1 and 2).
Under the null hypothesis, moderate adherence to the asymptotic properties shown in Theorem 3.3 is demonstrated. The NPGC permutation methodology tracks relatively close to the included reference CDFs in Figure 1 for TAR(2) simulation settings, and exhibits mild to moderate deviations for Lorenz-96 settings in Figure 2. These deviations can likely be attributed to the complicated nature of the chaotic Lorenz system; the same pattern is not observed in other simulations when both groups are generated by a TAR(2) process (see additional figures in the supplement). The cMLP lasso methods are included as horizontal lines in these plots; they do not have a straightforward extension for a specific level of test, other than computationally expensive iteration over several penalties $\lambda$. The naive methods do not adhere to their theoretical Type 1 error control.